OpenStudy (anonymous):

How do I set this up? In a basketball game, a player has scored 37 points on a total of 17 baskets consisting of 2-pointers and 3-pointers. How many of each basket has he scored?

OpenStudy (anonymous):

Let x = the number of 2 pointers. Let y = the number of 3 pointers So you have x+y=17 3x+2y=37

OpenStudy (anonymous):

Is it okay if I have 2x+3y=37 written down instead?

OpenStudy (anonymous):

You are correct. Sorry, I'm multitasking. x+y=17 2x+3y=37

OpenStudy (anonymous):

Would the answer be: 2-pointers:14 3-pointers: 3

OpenStudy (anonymous):

You got it.

OpenStudy (anonymous):

Thank you for your help :)